OpenStudy (anonymous):

for the given statement on Pn write the statements P1 Pk and Pk+1 2 + 4 + 6 + .... +2n = n(n+1)

OpenStudy (anonymous):

P1=1(1+1)=2 for Pk substitute n=k on both sides and then add 2(k+1) on both sides to get Pk+1

OpenStudy (anonymous):

@surjithayer so Pk would be what? I'm not sure what I plug in for it and where

OpenStudy (anonymous):

Pk is 2+4+6+...+2k=k(k+1)

OpenStudy (anonymous):

@surjithayer then do I solve for K?

OpenStudy (amistre64):

k is a general placeholder ... we know that for k=1, this holds if it holds for some k, then does it hold for k+1? is what we need to prove. but thats not what this is asking is it

OpenStudy (amistre64):

add 2(k+1) to both sides and simplify the right side to see if its of the form: (k+1) (k+1+1)

OpenStudy (anonymous):

So P1= 1(1+1)=2 Pk= 2+4+6+...+2k=k(k+1) Pk+1= 2+4+6+...+2(k+1)=k+1((k+1)+1) ?????